OpenStudy (anonymous):

What is the slope of the line 2x - 5y = -10? Answer none of the above

OpenStudy (anonymous):

this is a linear form, so what u want to do is isolate for your y first bring ur 2x to the other side giving u -5y=-2x-10 = 5y=2x+10 then just divide by 5 and u get y=2/5x+2 and the slope is a which is 2/5 :)

OpenStudy (anonymous):

jazy its not -